# OpenOOD v1.5: Enhanced Benchmark for Out-of-Distribution Detection

### v4

@@ Appendix C Related Work
Tajwar et al. [72] made the observation that “OOD detection methods are inconsistent across datasets” from experiments on 3 small datasets (CIFAR-10, CIFAR-100, and SVHN) with 3 specific post-hoc methods (MSP, ODIN, and MDS). While we draw a similar conclusion of “no single winner” in Section 5, our observation comes from the experimentation with 4 datasets and nearly 40 methods from different categories.
@@
A recent work by Galil et al. [73] proposed a method for constructing OOD detection benchmark and evaluated the performance of 5 post-hoc methods with ImageNet-1K pre-trained models. Specifically, for a specific ImageNet-1K model with a specific post-processor, they consider ImageNet-21K images as OOD and categorizes OOD images into a sequence of difficulty groups based on the OOD score from the post-processor. Correspondingly, their evaluation looks at the OOD detection AUROC across all groups, intending to provide a spectrum of AUROC v.s. difficulty. We see two shortcomings of such practice for constructing a general benchmark. First, their benchmarking process is extremely time-consuming since it needs to iterate through nearly all of the samples in ImageNet-21K, which could be prohibitive even for the most lightweight method considering the compute required by common ImageNet models. Second, the resulting benchmark is diagnostic to both the classifier and the post-processor. For example, the first difficulty group of the benchmark for MSP and that for ASH would not contain the same OOD samples, making the comparison ambiguous and much less straightforward. In comparison, our carefully designed benchmarks are standardized, i.e., agnostic to classifiers and post-processors. Plus, we consider a wide range of methods beyond a few specific post-hoc approaches.
@@
One work that most closely relates to ours is PyTorch-OOD [74], which is a python library for evaluating OOD detection performance. There are several distinctions that separate OpenOOD from PyTorch-OOD. 1) Number of supported methods. PyTorch-OOD implements 19 methods as of May 2023 with the most recent one dating back to 2022, while OpenOOD supports 40 approaches including the most advanced ones published in 2023. 2) Reliability of evaluation results. PyTorch-OOD still includes as OOD images the LSUN-R and TIN-R [5] which contain obvious resizing artifacts [7]. It also considers ImageNet-O which is known to cause biased evaluation since it is constructed by adversarially targeting a ResNet-50 model with the MSP detector [73]. Their benchmarking results thus can be problematic and unreliable. 3) Alignment between the goal reflected by the evaluation and human perception. PyTorch-OOD’s evaluation setup favors detectors that flag covariate-shifted ID samples (e.g., those from ImageNet-R or ImageNet-C) as OOD. We argue that this does not align with human perception and is not an ideal behavior as thoroughly discussed in Section 2 and Section 5.
@@
Another concurrent work by Bitterwolf et al. [68] put up a noise-free OOD dataset (NINCO) for ImageNet-1K in response to the observed noise that exists in popular OOD datasets. They then evaluate 8 post-hoc methods on NINCO and specifically study the effect of large-scale pre-training. OpenOOD is inherently complementary to the work of [68], as we intend to build a comprehensive benchmark for OOD detection by implementing and evaluating various types of methods (not restricting to post-hoc ones) on multiple datasets including ImageNet-1K. Meanwhile, our investigation on full-spectrum detection and findings regarding data augmentation techniques are unique.

### v5

@@ Appendix E Related Work
Tajwar et al. (2021) made the observation that “OOD detection methods are inconsistent across data sets” from experiments on 3 small data sets (CIFAR-10, CIFAR-100, and SVHN) with 3 specific post-hoc methods (MSP, ODIN, and MDS). While we draw a similar conclusion of “no single winner” in Section 6, our observation comes from the experimentation with 4 data sets and nearly 40 methods from different categories.
@@
A recent work by Galil et al. (2023) proposed a method for constructing OOD detection benchmark and evaluated the performance of 5 post-hoc methods with ImageNet-1K pre-trained models. Specifically, for a specific ImageNet-1K model with a specific post-processor, they consider ImageNet-21K images as OOD and categorizes OOD images into a sequence of difficulty groups based on the OOD score from the post-processor. Correspondingly, their evaluation looks at the OOD detection AUROC across all groups, intending to provide a spectrum of AUROC v.s. difficulty. We see two shortcomings of such practice for constructing a general benchmark. First, their benchmarking process is extremely time-consuming since it needs to iterate through nearly all of the samples in ImageNet-21K, which could be prohibitive even for the most lightweight method considering the compute required by common ImageNet models. Second, the resulting benchmark is diagnostic to both the classifier and the post-processor. For example, the first difficulty group of the benchmark for MSP and that for ASH would not contain the same OOD samples, making the comparison ambiguous and much less straightforward. In comparison, our carefully designed benchmarks are standardized, i.e., agnostic to classifiers and post-processors. Plus, we consider a wide range of methods beyond a few specific post-hoc approaches.
@@
One work that most closely relates to ours is PyTorch-OOD Kirchheim et al. (2022), which is a python library for evaluating OOD detection performance. There are several distinctions that separate OpenOOD from PyTorch-OOD. 1) Number of supported methods. PyTorch-OOD implements 19 methods as of May 2023 with the most recent one dating back to 2022, while OpenOOD supports 40 approaches including the most advanced ones published in 2023. 2) Reliability of evaluation results. PyTorch-OOD still includes as OOD images the LSUN-R and TIN-R Liang et al. (2018) which contain obvious resizing artifacts Tack et al. (2020). It also considers ImageNet-O which is known to cause biased evaluation since it is constructed by adversarially targeting a ResNet-50 model with the MSP detector Galil et al. (2023). Their benchmarking results thus can be problematic and unreliable. 3) Alignment between the goal reflected by the evaluation and human perception. PyTorch-OOD’s evaluation setup favors detectors that flag covariate-shifted ID samples (e.g., those from ImageNet-R or ImageNet-C) as OOD. We argue that this does not align with human perception and is not an ideal behavior as thoroughly discussed in Section 2 and Section 6.
@@
Another concurrent work by Bitterwolf et al. (2023) put up a new OOD data set (NINCO) for ImageNet-1K in response to the observed noise that exists in some earlier OOD data sets. They then evaluate 8 post-hoc methods on NINCO and specifically study the effect of large-scale pre-training. OpenOOD is inherently complementary to that work, as we intend to build a comprehensive benchmark for OOD detection by implementing and evaluating various types of methods (not restricting to post-hoc ones) on multiple data sets including ImageNet-1K. Meanwhile, our investigation on full-spectrum detection and findings regarding data augmentation techniques are unique.